# Urban Traffic Forecasting with Integrated Travel Time and Data Availability in a Conformal Graph Neural Network Framework

### v1

@@ Abstract
Traffic flow prediction is a big challenge for transportation authorities as it helps in planning and developing better infrastructure. State-of-the-art models often struggle to consider the data in the best way possible, intrinsic uncertainties, and the actual physics of the traffic. In this study, we propose a novel framework to incorporate travel times between stations into a weighted adjacency matrix of a Graph Neural Network (GNN) architecture with information from traffic stations based on their data availability. To handle uncertainty, we utilized the Adaptive Conformal Prediction (ACP) method that adjusts prediction intervals based on real-time validation residuals. To validate our results, we model a microscopic traffic scenario and perform a Monte-Carlo simulation to get a travel time distribution for a Vehicle Under Test (VUT) while it is navigating the traffic scenario, and this distribution is compared against the actual data. Experiments show that the proposed model outperformed the next best model by approximately 24% in MAE and 8% in RMSE and validation showed the simulated travel time closely matches the 95th percentile of the observed travel time value.
@@ I INTRODUCTION
Traffic congestion in cities is a big problem. They impact our daily commutes longer, increase pollution, and generally make urban life difficult. According to INRIX, the average time a U.S. person spent was 51 hours just sitting in traffic in 2022 — 15 hours more than in 2021 due to the pandemic but it was 99 hours in 2019 [1]. So, it is a time as well as a monetary loss. Knowing the traffic conditions a-priori can help in making better plans and avoiding congestion. Conventionally, studies have been conducted to predict traffic using historical data and statistical methods, which might not be able to capture the complex dynamics of a traffic scenario.
@@
Lately, there has been a lot of progress in using machine learning, especially in understanding patterns in a spatio-temporal, to get better at predicting how traffic will behave. By looking at roads like a network of nodes and edges (like a map) and using models that are good at handling sequences of data (like how traffic changes throughout the day), we can make smarter decisions. Specifically, using spatial models like Graph Neural Network (GNN) (which are appropriate for mapping out roads and intersections) with a model that can handle the time-varying nature of traffic such as a Recurrent Neural Network (RNN) can help tackle the complexities of forecasting traffic.
@@
However, even with advancements in the spatio-temporal domain, it is an arduous task to perfectly match these models with the actual physics, the associated uncertainties, and how the data is used in the best way possible to make traffic-related predictions. We leveraged Graph Convolution Network (GCN) with a Long Short-Term Memory (LSTM) architecture, inspired by Bing et al. [2], in this study. GCN is a type of GNN that applies the Convolutional Neural Network (CNN) concept in graph-structured data. It can perform a convolutional operation directly on the graph nodes and their neighboring nodes. The fundamental idea is to aggregate features of a selected node and its neighbors within an adjacency matrix. We propose to use travel times between nodes and their neighbors to make the graph model reflect a sense of realism. This helps us understand not just the layout of the roads but also how long it takes to get from one point to another. An LSTM architecture is used as they are known for their ability to capture long-term dependencies in sequential data which includes memory cells and gating mechanisms. This helps them to remember information over long sequences, addressing the vanishing gradient problem common in traditional RNNs.
We also look at the type of data coming from stations that may have a continuous count for the whole year with high granularity known as Continuous Count Stations (CCS) and some may be coarse known as Non-Continuous Count Stations (N-CCS). CCS provides detailed traffic counts for every 15/30/60 minutes for each day throughout the entire year whereas, N-CCS has limited data consisting of traffic counts for just a week or two, or scattered days, which, despite their sparsity, can still provide valuable insights for the traffic flow in that area. By using a weighting strategy to combine all this information, our model can make better predictions. Along with the predictions, uncertainty comes along so we used an adaptive version of conformal prediction known as ACP technique to deal with it. ACP is capable of providing dynamic prediction intervals or uncertainty bounds that adapt over time in a training model. Following this, a microscopic traffic scenario is modeled in SUMO (Simulation of Urban MObility) from the uncertain bounds of the predicted traffic flows. The validation is conducted through a Monte-Carlo simulation and analyzing the VUT travel time distribution against the actual data.
@@ II-A Spatio-Temporal Models
In 2019, the Temporal Graph Convolutional Network (T-GCN) was developed to improve the integration of spatial and temporal aspects in the traffic forecasts by combining graph convolutions with gated recurrent units [4]. In order to better capture dynamic traffic patterns, attention mechanisms were concurrently added by the Attention-based Spatial-Temporal Graph Convolutional Networks (ASTGCN) [5]. The Graph Multi-Attention Network (GMAN) and Graph WaveNet were also developed around this time and they improved the model in terms of managing intricate spatial-temporal interactions by utilising various attention processes [6], [7]. Bi-component graph convolutions and the integration of graph convolution into LSTM cells were introduced by other models such as the Temporal Graph Convolutional Long Short-Term Memory (TGC-LSTM) network and the Multi-range Attentive Bicomponent Graph Convolutional Network (MFFB) [8], [9]. We also saw the introduction of the Spatio-Temporal Meta learning framework (ST-MetaNet), applying meta-learning to spatio-temporal prediction, and the Dynamic Graph Convolutional Network (DGCN), which adjusted graph convolutional operations for constantly changing graphs [10], [11]. Then came Multi-Weighted Temporal Graph Convolutional network (MW-TGC) that refined spatial-temporal modeling even further by bringing in new ways to weigh temporal relationships and melding them with LSTM architectures [12]. Another study introduced the Structure Learning Convolution (SLC) framework [13] which bolstered teh traditional CNN architecture by incorporating structural information into the convolutional process and made it a versatile tool for learning on graphs.
@@
These advancements show how spatiotemporal models may improve urban traffic management and planning, as each has made a distinct contribution to the evolution of traffic flow prediction.
@@ II-B Physics-Based Deep Learning Models
While spatio-temporal models offer a significant improvement over prediction efficacies, the integration of physics-based principles into these models brings an additional layer of realism. Models like Greenshields model [16] or the more comprehensive first-order Lighthill-Whitham-Richards (LWR) model, gives a fundamental relationship between traffic density, flow, and speed [17] as described below.
@@
where $\rho$ represents traffic density, $v$ is the speed of the traffic, $x$ is the spatial coordinate, and $t$ is time. Also, higher-order models like Payne–Whitham(PW) model [18] and Aw–Rascle–Zhang(ARZ) models [19] were developed to overcome the limitations intrinsic to LWR by introducing additional variables and considerations. The PW model is given by equation 2 and 3.
$\frac{\partial\rho}{\partial t}+\frac{\partial(\rho v)}{\partial x}$ $=0,$ (2)
$\frac{\partial v}{\partial t}+v\frac{\partial v}{\partial x}$ $=-\frac{1}{\rho}\frac{\partial P}{\partial x}$ (3)
@@
$\frac{\partial\rho}{\partial t}+\frac{\partial(\rho v)}{\partial x}$ $=0,$ (4)
$\frac{\partial y}{\partial t}+v\frac{\partial y}{\partial x}$ $=0$ (5)
@@
where $v$ is the vehicle’s current speed, $v_{\text{desired}}$ is its desired speed, $\Delta v$ is the speed difference with the vehicle ahead, $s$ is the current gap, $s^{*}$ is the desired safety distance, $s_{0}$ is the minimum static gap, $T$ is the time headway, $a$ is the maximum acceleration, and $b$ is the comfortable braking deceleration. SUMO uses IDM to generate vehicle routes based on edge count data to simulate realistic vehicle interactions within the network. Studies like Physics-Informed Neural Networks (PINNs) utilized the LWR model in the presence of sparse data grids and proved their efficacy [20]. Additionally, Spatio-Temporal Dynamic Network (STDEN) model utilizes potential energy fields in order to take into account the dynamic impact of adjacent regions, consequently improving forecasting capabilities [21]. These research added an extra computational expense by incorporating physics-based parameters into the neural network training loss functions. Therefore, we suggest adding travel/transit times between nearby nodes to the adjacency matrix in order to address that. This intuitively makes sense as it is directly related to the fundamental parameters-speed, density, and flow.
@@ II-C Uncertainty Quantification
Predicting traffic flow accurately is challenging because there are so many unpredictable elements like sudden accidents, changing weather, and the number of cars on the road at any time. The traditional ways to predict the range of these predictions, use Bayesian inference, quantile regression, and other ensemble-based models [22], [23], and [24]. Apart from these ”parametric” methods, there is a comparatively newer technique called Conformal Prediction (CP) that has gotten a lot of attention as it comes with a coverage guarantee of the actual outcomes and it does not depend on a well-calibrated model. It also yields a transparent and traceable way of obtaining confidence intervals [25]. There is an even more advanced version of CP called Adaptive Conformal Prediction (ACP) [26], which changes the prediction ranges as new data comes in. It keeps updating these ranges to make sure they stay accurate even when things change and we modified [26] method to a simpler version by not considering an additional factor $\gamma$ (learning rate) but directly using the residuals (differences in the past prediction values) for maintaining the desired confidence level.
@@ III-A Graph Network
The road network is represented as a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ denotes the set of nodes that are traffic count stations and $\mathcal{E}$ represents the set of edges that connects these nodes. Each node $v_{i}\in\mathcal{V}$ has traffic flow measurements, and each edge $e_{ij}\in\mathcal{E}$ shows the road stretch between two nodes $v_{i}$ and $v_{j}$. The adjacency matrix $\mathbf{A}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$ is an important component in a graph network, that contains the information regarding connectivity and interaction between nodes. We introduced $\mathbf{A}$ based on travel times derived from average speeds and distances between stations, alongside data availability scores for CCS and N-CCS stations. Let us assume, $C_{i}$ denotes the number of counts at N-CCSi. The data availability for N-CCS stations can formulated by score normalization as:
$A_{i}^{NCCS}=\frac{C_{i}}{\max_{j}(C_{j})}$ (8)
where $C_{i}$ denotes counts at station $NCCS_{i}$ and $\max_{j}(C_{j})$ is the maximum count across all N-CCSs. The CCS stations are assumed fully available, denoted by $\mathbf{A}^{\text{CCS}}=\mathbf{1}_{|\mathcal{V}_{\text{CCS}}|}$, where $\mathcal{V}_{\text{CCS}}$ is the set of CCS nodes. The travel times $\mathbf{T}$, are computed as:
@@ III-B Graph Convolution Network (GCN)
In our model, we used a graph convolution network similar to [2] to understand spatial interconnections of the road network. In a GCN, nodes represent traffic stations, and edges signify the connecting roads. Each node gets a set of features and also information about its neighboring nodes. We update this combined information in steps. In each step, every point looks at the information of the points it is directly connected to and gathers up this information known as a message-passing mechanism:
@@
where $\mathbf{h}_{i}^{(l)}$ denotes the information of node $i$ at layer $l$, $\mathcal{N}(i)$ represents the set of neighbors of node $i$, $\text{A}^{(l)}(\cdot)$ is an aggregation function that combines the information of neighboring nodes, and $\text{U}^{(l)}(\cdot)$ is an update function that generates the new information for node $i$ based on its previous embedding and the aggregated message. The network then updates the node embeddings using a combination of the aggregated messages and the nodes previous embeddings, often through a transformation that involves learnable parameters and non-linear activation functions.
@@ III-C LSTM with Attention Mechanism
To understand how traffic patterns change temporally, we used an LSTM model as it is suited for capturing patterns in sequential manner, like how busy a road gets during the day. LSTM looks at the information from the GCN in sequence. However, we wanted our model to pay more attention to the most important moments in the data. For this, we leveraged the attention mechanism [27]. It assigns a weight, or importance level, to each relevant time step based on how much it should influence the final prediction:
@@ III-D Adaptive Conformal Prediction (ACP)
We used a modified CP framework called ACP to quantify the uncertainty in the predictions. It provides prediction intervals that encompass the actual traffic flow values. To maintain the accuracy, it modifies the intervals in response to the observed residuals [26]. It then calculates the residuals between the actual and predicted values and then defines the quantile value from these residuals.
@@
where $q_{\text{adjusted}}$ is the quantile of the absolute residuals observed on the validation set, corresponding to a significance level $\alpha$ that represents the proportion of future predictions we expect to fall inside the prediction intervals. The adaptive mechanism updates $q_{\text{adjusted}}$ at the end of each training epoch based on the latest residuals. A callback function then evaluates the performance on the validation set, recalculates $q_{\text{adjusted}}$, and adjusts the prediction intervals accordingly. We used $\alpha=0.1$, meaning we expected $90\%$ of the times the true values will fall within the predicted interval.
@@ III-E Optimization
In training the model, we used Mean Squared Error (MSE) as the loss function in our model.
@@
where $N$ is the number of samples, $\hat{y}_{i}$ is the predicted traffic flow, and $y_{i}$ is the actual traffic flow. To help our model learn better and faster, we used RMSprop optimization algorithm with a learning rate of 0.0002. To avoid overfitting, we used an early stopping technique that stops the training, if the model does not show improvement over a certain epoch. Also, we used a custom callback function to update the quantile used in ACP based on the residuals observed on the validation dataset. This keeps our predictions in check, making sure that the ranges we predict ensure how well our model is doing at any point in time.

### v2

@@ Abstract
Traffic flow prediction is a big challenge for transportation authorities as it helps plan and develop better infrastructure. State-of-the-art models often struggle to consider the data in the best way possible, as well as intrinsic uncertainties and the actual physics of the traffic. In this study, we propose a novel framework to incorporate travel times between stations into a weighted adjacency matrix of a Graph Neural Network (GNN) architecture with information from traffic stations based on their data availability. To handle uncertainty, we utilized the Adaptive Conformal Prediction (ACP) method that adjusts prediction intervals based on real-time validation residuals. To validate our results, we model a microscopic traffic scenario and perform a Monte-Carlo simulation to get a travel time distribution for a Vehicle Under Test (VUT), and this distribution is compared against the real-world data. Experiments show that the proposed model outperformed the next-best model by approximately 24% in MAE and 8% in RMSE and validation showed the simulated travel time closely matches the 95th percentile of the observed travel time value.
@@ I INTRODUCTION
Traffic congestion in cities significantly impacts daily commutes, increases pollution, and makes urban life challenging. According to INRIX [1], the average U.S. commuter spent 51 hours in traffic in 2022 — 15 hours more than in 2021 down from 99 hours in 2019. This persistent issue underscores the need for accurate traffic prediction. Predicting traffic conditions can aid in planning and avoiding congestion, yet traditional methods using historical data often fail to capture the complex dynamics of traffic effectively.
@@
Lately, there has been significant progress in using machine learning techniques, particularly in understanding patterns in a spatio-temporal domain to predict traffic behavior. By viewing roads as a network of nodes and edges similar to a map, and utilizing models adept at handling sequential data such as how traffic changes throughout the day, we can make more informed decisions. Specifically, employing spatial models like Graph Neural Networks (GNNs) which are well-suited for mapping roads and intersections alongside models that handle the time-varying nature of traffic such as Recurrent Neural Networks (RNNs) can help address the complexities of traffic forecasting.
@@
However, even with advancements in the models, it is still an arduous task to account for the actual physics, the associated uncertainties, and how the data is used in the best way possible. We leveraged Graph Convolution Network (GCN) with a Long Short-Term Memory (LSTM) architecture, inspired by Bing et al. [2], in this study. GCN is a type of GNN that applies the Convolutional Neural Network (CNN) concept in graph-structured data. It can perform a convolutional operation directly on the graph nodes and their neighboring nodes. Fundamentally, it can aggregate features of a selected node and its neighbors within an adjacency matrix. We propose to use travel times between nodes and their neighbors (traffic stations) to make the graph model reflect a sense of realism. This helps us understand not just the layout of the roads but also how long it takes to get from one point to another. An LSTM architecture captures long-term dependencies in sequential data, using memory cells and gating mechanisms. This helps them to remember information over long sequences, addressing the vanishing gradient problem common in traditional RNNs.
We utilize data from Continuous Count Stations (CCS) and Non-Continuous Count Stations (N-CCS), combining high granularity data with sparse but insightful data. CCS provides detailed traffic counts every 15, 30, or 60 minutes year-round, while N-CCS offers limited data for short periods, adding sporadic yet valuable insights. To enhance predictions, we employ a weighting strategy that effectively balances the detailed CCS data with the sparse N-CCS data. Additionally, an adaptive conformal prediction (ACP) technique is employed to quantify uncertainty. To verify the effectiveness of our approach, validation is conducted by modeling a microscopic traffic scenario in SUMO (Simulation of Urban MObility) using predicted traffic flows and analyzing the VUT travel time distribution through Monte-Carlo simulation against real-world INRIX data.
@@ II-A Spatio-Temporal Models
In 2019, the Temporal Graph Convolutional Network (T-GCN) was developed to improve the integration of spatial and temporal aspects in traffic forecasts by combining graph convolutions with gated recurrent units [4]. In order to better capture dynamic traffic patterns, attention mechanisms were concurrently added by the Attention-based Spatial-Temporal Graph Convolutional Networks (ASTGCN) [5]. The Graph Multi-Attention Network (GMAN) and Graph WaveNet were also developed around this time and they improved the model in terms of managing intricate spatial-temporal interactions by utilizing various attention processes [6], [7]. Bi-component graph convolutions and the integration of graph convolution into LSTM cells were introduced by other models such as the Temporal Graph Convolutional Long Short-Term Memory (TGC-LSTM) network and the Multi-range Attentive Bicomponent Graph Convolutional Network (MFFB) [8], [9]. We also saw the introduction of the Spatio-Temporal Meta-learning framework (ST-MetaNet), applying meta-learning to spatio-temporal prediction, and the Dynamic Graph Convolutional Network (DGCN), which adjusted graph convolutional operations for constantly changing graphs [10], [11]. Then came the Multi-Weighted Temporal Graph Convolutional network (MW-TGC) that refined spatial-temporal modeling even further by bringing in new ways to weigh temporal relationships and melding them with LSTM architectures [12]. Another study introduced the Structure Learning Convolution (SLC) framework [13] which bolstered the traditional CNN by incorporating structural information into the convolutional process.
@@
These advancements show how spatio-temporal models may improve urban traffic management and planning, as each has made a distinct contribution to the evolution of traffic state prediction.
@@ II-B Physics-Based Deep Learning Models
While spatio-temporal models improve prediction efficacy, integrating physics-based principles adds realism. These models can capture fundamental relationships between traffic density, flow, and speed, reflecting the physical dynamics of traffic. Models like the Greenshields model [16] or the first-order Lighthill-Whitham-Richards (LWR) model provide insights into these interactions [17], as shown in equation 1.
@@
where $\rho$ represents traffic density, $v$ is the speed of the traffic, $x$ is the spatial coordinate, and $t$ is time. Also, higher-order models like Payne–Whitham (PW) model [18] and Aw–Rascle–Zhang (ARZ) models [19] were developed to overcome the limitations intrinsic to LWR by introducing additional variables and considerations. The PW model is given by equations 2 and 3.
$\frac{\partial\rho}{\partial t}+\frac{\partial(\rho v)}{\partial x}$ $=0,$ $\frac{\partial v}{\partial t}+v\frac{\partial v}{\partial x}$ $=-\frac{1}{\rho}\frac{\partial P}{\partial x}$ (2)
@@
$\frac{\partial\rho}{\partial t}+\frac{\partial(\rho v)}{\partial x}$ $=0,$ $\frac{\partial y}{\partial t}+v\frac{\partial y}{\partial x}$ $=0$ (3)
@@
where $v$ is the vehicle’s current speed, $v_{\text{desired}}$ is its desired speed, $\Delta v$ is the speed difference with the vehicle ahead, $s$ is the current gap, $s^{*}$ is the desired safety distance, $s_{0}$ is the minimum static gap, $T$ is the time headway, $a$ is the maximum acceleration, and $b$ is the comfortable braking deceleration. SUMO uses IDM to generate vehicle routes based on edge count data to simulate realistic vehicle interactions within the network. Later studies like Physics-Informed Neural Networks (PINNs) utilized the LWR model in the presence of sparse data grids and proved their efficacy [20]. Additionally, the Spatio-Temporal Dynamic Network (STDEN) model utilized potential energy fields in order to take into account the dynamic impact of adjacent regions, consequently improving forecasting capabilities [21].
While the above-discussed investigations improve accuracy, they introduced an extra computational expense by incorporating physics-based parameters into the neural network training loss functions. To address this, we propose adding the travel/transit times between the traffic stations to the adjacency matrix to reduce computational overhead while maintaining accuracy. This intuitively makes sense as it is directly related to the fundamental parameters- speed, density, and flow.
@@ II-C Uncertainty Quantification
Traffic flow prediction comes with many unpredictable elements like sudden accidents, changing weather, and the number of cars on the road at any time. The traditional ways to quantify uncertainties use Bayesian inference, quantile regression, and other ensemble-based models [22], [23], and [24]. There is a comparatively newer technique called Conformal Prediction (CP) that has gotten a lot of attention as it comes with a coverage guarantee of the actual outcomes and it does not depend on a well-calibrated model. It also yields a transparent and traceable way of obtaining confidence intervals [25]. There is an even more advanced version of CP called Adaptive Conformal Prediction (ACP) [26], which changes the prediction ranges as new data comes in. It keeps on updating these ranges to make sure they stay accurate even when the incoming data changes. However, we modified this to a simpler form by not considering the additional factor $\gamma$ (learning rate) and directly used the residuals (differences in the past prediction values) to maintain the desired confidence level.
@@ III-A Graph Network
The road network is represented as a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ denotes the set of nodes that are traffic count stations and $\mathcal{E}$ represents the set of edges that connects these nodes. Each node $v_{i}\in\mathcal{V}$ has traffic flow measurements, and each edge $e_{ij}\in\mathcal{E}$ shows the road stretch between two nodes $v_{i}$ and $v_{j}$. Here, $i$ and $j$ are the indices representing nodes in the traffic network. The adjacency matrix $\mathbf{A}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$ is an important component in a graph network, that contains the information regarding connectivity and interaction between nodes. We introduced $\mathbf{A}$ based on travel times derived from average speeds and distances between the traffic stations, alongside data availability scores for CCS and N-CCS stations. Let us assume, $C_{i}$ denotes the number of counts at N-CCSi. The data availability for N-CCS stations can formulated by score normalization as:
$A_{i}^{N-CCS}=\frac{C_{i}}{\max_{j}(C_{j})}$ (6)
where $C_{i}$ denotes counts at station $N-CCS_{i}$ and $\max_{j}(C_{j})$ is the maximum count across all N-CCSs. The CCS stations are assumed fully available, denoted by $\mathbf{A}^{\text{CCS}}=\mathbf{1}_{|\mathcal{V}_{\text{CCS}}|}$, where $\mathcal{V}_{\text{CCS}}$ is the set of CCS nodes. The travel times $\mathbf{T}$, are computed as:
@@ III-B Graph Convolution Network (GCN)
In our model, we used GCN [2] to understand spatial interconnections of the road network. In a GCN, nodes represent traffic stations, and edges signify the connecting roads. Each node gets a set of features and also information about its neighboring nodes. We update this combined information in steps. In each step, every point looks at the information of the points it is directly connected to and gathers up this information known as a message-passing mechanism:
@@
where ${m}_{i}^{(l+1)}$ represents the aggregated message for node $i$ at layer $l+1$, $\text{A}^{(l)}(\cdot)$ combines the information of neighboring nodes ${h}_{j}^{(l)}$, $j$ is in the set of neighbors $N(i)$ of node $i$, ${h}_{i}^{(l)}$ denotes the information of node $i$ at layer $l$, and $\text{U}^{(l)}(\cdot)$ is an update function generating new information for node $i$ based on its previous embedding and the aggregated message. The network then updates the node embeddings using a combination of the aggregated messages and the nodes’ previous embeddings, often through a transformation that involves learnable parameters and non-linear activation functions.
@@ III-C LSTM with Attention Mechanism
We used an LSTM model to capture temporal traffic patterns such as daily road congestion. LSTM processes GCN data sequentially but to emphasize on important moments, we added an attention mechanism [27]. This mechanism assigns weights to each time step, enhancing the final prediction by focusing on the most critical data points.
@@ III-D Adaptive Conformal Prediction (ACP)
We used a modified CP framework called ACP to quantify the uncertainty in the predictions. ACP provides prediction intervals that encompass actual traffic flow values and adjusts these intervals based on observed residuals [26]. Residuals between actual and predicted values are calculated, defining the quantile value:
@@
where $\hat{F}$ is the predicted traffic flow, and $q_{\text{adjusted}}$ is the quantile of the absolute residuals observed on the validation set, corresponding to a significance level $\alpha$ that represents the proportion of future predictions we expect to fall inside the prediction intervals. The adaptive mechanism updates $q_{\text{adjusted}}$ at the end of each training epoch based on the latest residuals. A callback function then evaluates the performance on the validation set, recalculates $q_{\text{adjusted}}$, and adjusts the prediction intervals accordingly. We used $\alpha=0.1$, meaning we expected $90\%$ of the times the true values will fall within the predicted interval.
@@ III-E Optimization
In training the model, we used Mean Squared Error (MSE) as the loss function.
@@
where $N$ is the number of samples, $\hat{y}_{i}$ is the predicted traffic flow, and $y_{i}$ is the actual traffic flow. We used the RMSprop optimization algorithm with a learning rate of 0.0002 and is dynamically adjustable for each parameter to improve convergence speed and handle non-stationary objectives. To prevent overfitting, we implemented an early stopping technique that halts training if no improvement is seen over a certain epoch. Additionally, a custom callback function updates the ACP quantile based on validation dataset residuals ensuring accurate prediction ranges.
The overall model architecture showing interactions between its components (GCN layer for spatial data, an LSTM with Attention for temporal patterns, and an ACP layer for uncertainty quantification) is illustrated in Figure 1.
Figure 1: Overall Model Architecture
@@
The techniques employed i.e. batch processing, early stopping, and RMSprop optimization, ensure computational efficiency and scalability. It also expedites training and enables the effective handling of large datasets, making the system suitable for real-time traffic prediction.